Grade 12Algebra

Durgesh Kushwaha
13 Years agoGrade 12

1 Answer

Aman  Bansal
13 Years ago

Dear Durgesh,

Express as a partial fraction


	 x - 1
    ---------------
    (3x - 5)(x - 3)


========================

Solution:

First write the fraction as

	  x - 1               A                B
     ---------------   =   --------    +    -------
     (3x - 5)(x - 3)       (3x - 5)         (x - 3)


Notice that I have taken the two terms that are in brackets and placed them on their own
as seperate fractions using ''A'' and ''B'' as numerators. 

A and B are the mystery numbers we need to discover.

Next, multiply denominator (the bottom bit of the big fraction) by both sides of the equation to cancel out terms



So we now have

	  x - 1       =   A(x - 3)   +   B(3x - 5)


In order to get rid of either the ''A'' or ''B'' term, we can substitute a ''strategic'' value to make it equal to zero. 
For example, to get rid of A(x - 3), let''s make x = 3, so (3 - 3) ends up as 0. Gone.

Remember that you have to do this to all of the x''s in the equation

	  3 - 1	      =   A(3 - 3)  +  B(3*3 - 5) you can see the ''A'' component becomes zero.

	  3 - 1	      =   B(9 - 5)

	  2	      =   4B

   so,	 2/4          =   B

   or,	 1/2          =   B


We''ve discovered what B is, now we can repeat the process to discover A

Let''s make x = 5/3 to get rid of the ''B'' part of the equation...

so,          B(5/3 * 3) = 5

and,         B(5 - 5)   = 0

The B(3x - 5) term is gone! So we can plug this value into the entire equation to find A


	  5/3 - 1      =   A(5/3 - 3)

	  2/3          =   A * -4/3

	  2/3
         ------        =   A
          -4/3

	  -1/2         =   A



We now have A and B, and the answer to our problem.


	 x - 1               -1              1
    ---------------   =   ---------   +   --------
    (3x - 5)(x - 3)       2(3x - 5)       2(x - 3)

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. 

https://www.askiitians.com/packages/packages.aspx

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert